Five hundred depth readings for an inland lake were normally distributed with a mean of 62 feet and a standard deviation of 8.5 feet. What percent of the depths were between 53.5 feet and 70.5 feet?
Answer by tommyt3rd(5050)

z1=(53.5-62)/8.5=-1
z2=(70.5-62)/8.5=1
P(-1 < z < 1)=0.6826
about 68%